OpenStudy (anonymous):

the probability that a professional baseball player will get a hit is 1/3. calculate the exact probability that he will get at least 3 hits in 5 attempts

OpenStudy (kropot72):

The binomial distribution is used for this problem. \[P(hits=3)=\left(\begin{matrix}5 \\ 3\end{matrix}\right)(\frac{1}{3})^{3}(\frac{2}{3})^{2}=\frac{40}{243}\]